# Provable Ordering and Continuity in Vision-Language Pretraining for Generalizable Embodied Agents

### v1

@@ Appendix A Proofs
For the proof of Theorem 4.2, we closely follow the approaches presented in  (Zha et al., 2023) and adapted to our triplet case. We prove the theorem in three steps:
@@
(3) For any $10<\delta<$ 1, there exist $\epsilon>0$, such that if $\mathcal{L}_{\mathrm{VLO}}<\mathcal{L}^{*}+\epsilon$, then the learned representations satisfy VLO property.
@@
$\mathcal{L}_{\mathrm{VLO}}=-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{j\in[T]\backslash\{i\}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}\geq d_{i,j}}\exp\left(\mathfrak{R}_{i,k,l}\right)}$ (7)
$=-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}\geq D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}$
$=-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{1}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}\geq D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}$
$=-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{1}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}$
$-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}\geq D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}$
$=-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}$
$+\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\left(1+\frac{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}>D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}\right)$
$>-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}.$
@@
$-\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}$ (8)
$\geq-n_{i,m}\log\left(\frac{1}{n_{i,m}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}\right)=n_{i,m}\log n_{i,m}.$
@@
$\left\{\begin{array}[]{l}\mathfrak{R}_{i,j,l}>\mathfrak{R}_{i,k,l}+\gamma\text{ if }d_{i,j}<d_{i,k}\\ \mathfrak{R}_{i,j,l}=\mathfrak{R}_{i,k,l}\text{ if }d_{i,j}=d_{i,k}\end{array}\right.$
and $\gamma:=\log\frac{T}{\min\limits_{i\in[T],m\in\left[M_{i}\right]}n_{i,m}\epsilon},\forall i\in[T],j,k\in[T]\backslash\{i\}$, such that $\mathcal{L}_{\mathrm{VLO}}<L^{\star}+\epsilon$. For such a set of representations, $\forall i\in[T],m\in\left[M_{i}\right],j\in\left\{[T]\backslash\{i\}\mid d_{i,j}=D_{i,m}\right\}$,
$-\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}=\log n_{i,m}$ (10)
@@
$\begin{array}[]{rl}&\log\left(1+\frac{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}>D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}\right)\\ &<\log\left(1+\frac{T\exp(-\gamma)}{n_{i,m}}\right)<\frac{T\exp(-\gamma)}{n_{i,m}}\leq\epsilon.\end{array}$ (11)
@@
$\mathcal{L}_{\mathrm{VLO}}=$ $-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}$ (12)
$+\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\left(1+\frac{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}>D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}\right).$
@@
$\epsilon=\frac{1}{T(T-1)}\min\left(\min_{i\in[T],m\in\left[M_{i}\right]}\log\left(1+\frac{1}{n_{i,m}\exp\left(\delta+\frac{1}{\delta}\right)}\right),2\log\frac{1+\exp(\delta)}{2}-\delta\right)>0,$
such that when $\mathcal{L}_{\mathrm{VLO}}<L^{*}+\epsilon$, the representations satisfy VLO property. We first show that $\left|\mathfrak{R}_{i,j,l}-\mathfrak{R}_{i,k,l}\right|<\delta$ if $d_{i,j}=d_{i,k}$, i $\in[T],j,k\in[T]\backslash\{i\}$ when $L_{\mathrm{VLO}}<L^{*}+\epsilon$. From Eq. (7) we have
$\mathcal{L}_{\operatorname{VLO}}>-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\sum\limits_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}-D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}$ (14)
Let $p_{i,m}:=\underset{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}{\arg\min}\mathfrak{R}_{i,j,l},q_{i,m}:=\underset{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}{\arg\max}\mathfrak{R}_{i,j,l},\zeta_{i,m}:=\mathfrak{R}_{i,p_{i,m},l},\eta_{i,m}:=$ $s_{i,q_{i,m},l}-s_{i,p_{i,m},l},\forall i\in[T],m\in\left[M_{i}\right]$, by splitting out the maximum term and the minimum term we have
$\mathcal{L}_{\mathrm{VLO}}>-\frac{1}{T(T-1)}\sum\limits_{i=1}^{T}\sum\limits_{m=1}^{M_{i}}\left\{\log\frac{\exp\left(\zeta_{i,m}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,i}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}\right.$ (15)
$\left.+\log\frac{\exp\left(\zeta_{i,m}+\eta_{i,m}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}+\log\frac{\exp\left(\sum\limits_{j\in[T]\backslash\left\{i,p_{i,m},q_{i,m}\right\},d_{i,j}=D_{i,m}}\mathfrak{R}_{i,j,l}\right)}{\left(\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)\right)^{n_{i,m}-2}}\right\}.$
Let $\theta_{i,m}:=\frac{1}{n_{i,m}-2}\sum\limits_{j\in[T]\backslash\left\{i,p_{i,m},q_{i,m}\right\},d_{i,j}=D_{i,j}}\exp\left(\mathfrak{R}_{i,j,l}-\zeta_{i,m}\right)$, we have
$-\log\frac{\exp\left(\zeta_{i,m}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}=\log\left(1+\exp\left(\eta_{i,m}\right)+\left(n_{i,m}-2\right)\theta_{i,m}\right)$ (16)
@@
$-\log\frac{\exp\left(\zeta_{i,m}+\eta_{i,m}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}=\log\left(1+\exp\left(\eta_{i,m}\right)+\left(n_{i,m}-2\right)\theta_{i,m}\right)-\eta_{i,m}$ (17)
@@
$\exp\left(\sum_{j\in[T]\backslash\left\{i,p_{i,m},q_{i,m}\right\},d_{i,j}=D_{i,m}}\mathfrak{R}_{i,j,l}\right)\leq\left(\frac{1}{n_{i,m}-2}\sum_{j\in[T]\backslash\left\{i,p_{i,m},q_{i,m}\right\},d_{i,j}=D_{i,m}}\exp\left(\mathfrak{R}_{i,j,l}\right)\right)^{n_{i,m}-2},$ (18)
@@
$-\log\frac{\exp\left(\sum\limits_{j\in[T]\backslash\{i,p_{i,m},q_{i,m}\},d_{i,j}=D_{i,m}}\mathfrak{R}_{i,j,l}\right)}{\left(\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)\right)^{n_{i,m}-2}}\geq\left(n_{i,m}-2\right)\log\left(1+\exp\left(\eta_{i,m}\right)+\left(n_{i,m}-2\right)\theta_{i,m}\right)-\left(n_{i,m}-2\right)\log\left(\theta_{i,m}\right)$ (19)
@@
$\mathcal{L}_{\mathrm{VLO}}>\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\left(n_{i,m}\log\left(1+\exp\left(\eta_{i,m}\right)+\left(n_{i,m}-2\right)\theta_{i,m}\right)-\eta_{i,m}-\left(n_{i,m}-2\right)\log\left(\theta_{i,m}\right)\right).$ (20)
@@
$\mathcal{L}_{\mathrm{VLO}}=$ $-\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\sum_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}\right)}$ (25)
$+\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\sum_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\left(1+\frac{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}>D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}\right),$
@@
$\mathcal{L}_{\mathrm{VLO}}$ $\geq L^{\star}+\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\sum_{j\in[T]\backslash\{i\},d_{i,j}=D_{i,m}}\log\left(1+\frac{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}>D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{k\in[T]\backslash\{i\},d_{i,k}=D_{i,m}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}\right)$ (26)
$>L^{\star}+\frac{1}{T(T-1)}\log\left(1+\frac{\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{h\in[T]\backslash\{i\},d_{i,h}=d_{i,j}}\exp\left(\mathfrak{R}_{i,h,l}-\mathfrak{R}_{i,j,l}\right)}\right),$
$\forall i\in[T],j\in[T]\backslash\{i\},k\in\left\{k\in[T]\backslash\{i\}\mid d_{i,j}<d_{i,k}\right\}$. When $\mathcal{L}_{\mathrm{VLO}}<L^{\star}+\epsilon$, we already have $\left|\mathfrak{R}_{i,h,l}-\mathfrak{R}_{i,j,l}\right|<\delta,\forall d_{i,h}=d_{i,j}$, which derives $\mathfrak{R}_{i,h,l}-\mathfrak{R}_{i,j,l}<\delta$ and thus $\exp\left(\mathfrak{R}_{i,h,l}-\mathfrak{R}_{i,j,l}\right)<\exp(\delta)$. By putting this into Eq. (25), we have $\forall i\in[T],j\in$ $[T]\backslash\{i\},k\in\left\{k\in[T]\backslash\{i\}\mid d_{i,j}<d_{i,k}\right\}$,

### v3

@@ C.1 Proofs of Theorem 1
For the proof of Theorem 1, we closely follow the approaches presented in [nips23-rnc] and adapted to our triplet case. We prove the theorem in three steps:
@@
(3) For any $0<\delta<1$, there exist $\epsilon>0$, such that if $\mathcal{L}_{\mathrm{VLO}}<\mathcal{L}^{*}+\epsilon$, then the learned representations satisfy VLO property.
@@
[…]
@@
[…]
@@
{R_i, j,l>R_i, k,l+γ if d_i, j<d_i, kR_i, j,l=R_i, k,l if d_i, j=d_i, k
[…]
@@
[…]
@@
$\mathcal{L}_{\mathrm{VLO}}=$ $-\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\sum_{\begin{subarray}{c}j\in[T]\setminus\{i\}\\ d_{i,j}=D_{i,m}\end{subarray}}\log\frac{\exp\left(\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}=D_{i,m}\end{subarray}}\exp\left(\mathfrak{R}_{i,k,l}\right)}$ (11)
$+\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\sum_{\begin{subarray}{c}j\in[T]\setminus\{i\}\\ d_{i,j}=D_{i,m}\end{subarray}}\log\left(1+\frac{\sum\limits_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}>D_{i,m}\end{subarray}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}{\sum\limits_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}=D_{i,m}\end{subarray}}\exp\left(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l}\right)}\right)$
@@
ϵ=1T(T-1) min(min_i ∈[T], m ∈[M_i] log(1+1ni, mexp(δ+1δ)), 2 log1+exp(δ)2-δ)>0,
[…]
@@
[…]
@@
$\exp\bigg(\sum_{\begin{subarray}{c}j\in[T]\setminus\{i,p_{i,m},q_{i,m}\}\\ d_{i,j}=D_{i,m}\end{subarray}}\mathfrak{R}_{i,j,l}\bigg)\leq\left(\frac{1}{n_{i,m}-2}\sum_{\begin{subarray}{c}j\in[T]\setminus\{i,p_{i,m},q_{i,m}\}\\ d_{i,j}=D_{i,m}\end{subarray}}\exp(\mathfrak{R}_{i,j,l})\right)^{n_{i,m}-2}$ (17)
@@
$-\log\frac{\exp\left(\sum_{\begin{subarray}{c}j\in[T]\setminus\{i,p_{i,m},q_{i,m}\}\\ d_{i,j}=D_{i,m}\end{subarray}}\mathfrak{R}_{i,j,l}\right)}{\left(\sum_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}=D_{i,m}\end{subarray}}\exp(\mathfrak{R}_{i,k,l})\right)^{n_{i,m}-2}}$ (18)
$\quad\geq(n_{i,m}-2)\log\left(1+\exp(\eta_{i,m})+(n_{i,m}-2)\theta_{i,m}\right)-(n_{i,m}-2)\log(\theta_{i,m})$
@@
$\mathcal{L}_{\mathrm{VLO}}>\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\bigg($ $n_{i,m}\log\left(1+\exp(\eta_{i,m})+(n_{i,m}-2)\theta_{i,m}\right)$ (19)
$-\eta_{i,m}-(n_{i,m}-2)\log(\theta_{i,m})\bigg)$
@@
$\mathcal{L}_{\mathrm{VLO}}=$ $-\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\sum_{\begin{subarray}{c}j\in[T]\setminus\{i\}\\ d_{i,j}=D_{i,m}\end{subarray}}\log\frac{\exp(\mathfrak{R}_{i,j,l})}{\sum\limits_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}=D_{i,m}\end{subarray}}\exp(\mathfrak{R}_{i,k,l})}$ (24)
$+\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\sum_{\begin{subarray}{c}j\in[T]\setminus\{i\}\\ d_{i,j}=D_{i,m}\end{subarray}}\log\left(1+\frac{\sum\limits_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}=D_{i,m}\end{subarray}}\exp(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l})}{\sum\limits_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}>D_{i,m}\end{subarray}}\exp(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l})}\right)$
@@
$\mathcal{L}_{\mathrm{VLO}}\geq$ $L^{\star}+\frac{1}{T(T-1)}\sum_{i=1}^{T}\sum_{m=1}^{M_{i}}\sum_{\begin{subarray}{c}j\in[T]\setminus\{i\}\\ d_{i,j}=D_{i,m}\end{subarray}}\log\left(1+\frac{\sum\limits_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}>D_{i,m}\end{subarray}}\exp(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l})}{\sum\limits_{\begin{subarray}{c}k\in[T]\setminus\{i\}\\ d_{i,k}=D_{i,m}\end{subarray}}\exp(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l})}\right)$ (25)
$>$ $\ L^{\star}+\frac{1}{T(T-1)}\log\left(1+\frac{\exp(\mathfrak{R}_{i,k,l}-\mathfrak{R}_{i,j,l})}{\sum\limits_{\begin{subarray}{c}h\in[T]\setminus\{i\}\\ d_{i,h}=d_{i,j}\end{subarray}}\exp(\mathfrak{R}_{i,h,l}-\mathfrak{R}_{i,j,l})}\right)$
$\forall i\in[T],j\in[T]\backslash\{i\},k\in\left\{k\in[T]\backslash\{i\}\mid d_{i,j}<d_{i,k}\right\}$. When $\mathcal{L}_{\mathrm{VLO}}<L^{\star}+\epsilon$, we already have $\left|\mathfrak{R}_{i,h,l}-\mathfrak{R}_{i,j,l}\right|<\delta,\forall d_{i,h}=d_{i,j}$, which derives $\mathfrak{R}_{i,h,l}-\mathfrak{R}_{i,j,l}<\delta$ and thus $\exp\left(\mathfrak{R}_{i,h,l}-\mathfrak{R}_{i,j,l}\right)<\exp(\delta)$. By putting this into Eq. (24), we have $\forall i\in[T],j\in$ $T]\backslash\{i\},k\in\left\{k\in[T]\backslash\{i\}\mid d_{i,j}<d_{i,k}\right\}$,